OpenStudy (anonymous):

Hannah wants to encircle a tree trunk with a flower bed. If the center of the tree trunk is the origin and Hannah wants the flower bed to extend 3 feet from the center of the tree, what is the equation that would represent the flower bed?

OpenStudy (anonymous):

xy=C

OpenStudy (anonymous):

not too sure but circumfrance is c and one of your points would be 3

OpenStudy (anonymous):

its either x2-y2=3 or x2+y2=3

OpenStudy (anonymous):

i would say second

OpenStudy (anonymous):

aha it was x2+y2=9. oh well, thanks tho(:

OpenStudy (anonymous):

yea cause adding 3 feet oh yea my math isnt all bad lol best resonse me please